# Improved particle-flow event reconstruction with scalable neural networks for current and future particle detectors

### 29 Jan 2024 v4

@@ 4 Conclusions
We have used a realistic simulation for the CLIC detector, conceptually similar to existing and future detector designs, to develop scalable machine learning models for full event particle flow reconstruction. Based on a large-scale hyperparameter optimization (HPO), we find that a graph neural network (GNN) model that uses locality sensitive hashing (LSH) significantly outperforms the baseline Pandora-based particle-flow (PF) on the basis of individual particles as well as event-level quantities such as jets and total 3-momentum. Improved particle-level and event-level reconstruction can result in significant improvements for future flagship analyses such as those involving the Higgs boson decaying to bottom quarks. At the same time, the flexible and learnable LSH approach allows to process complex events expected at future detectors with up to $10^{4}$ particles per event efficiently, while supporting various existing hardware accelerators such as Nvidia, AMD and Habana without additional porting effort. Our work contributes to the existing body of research by proposing a new, challenging open dataset for particle flow reconstruction studies, by defining relevant benchmarks and by identifying efficient models that can solve these benchmarks better than existing rule-based algorithms without additional tuning.
@@
Further research is possible in several directions. First, it would be useful to repeat this exercise using the simulation from detectors that are taking data in Run 3 of the Large Hadron Collider (LHC), e.g. in CMS, to study the performance of the reconstruction in more realistic conditions, and also study the reconstruction performance on real data. Second, we are currently using a simple particle-based loss, while the use of contrastive-adversarial learning methods may allow to account for event-level discrepancies more effectively [65]. Third, the hypergraph model [21] shows promising physics performance but currently only supports small input sequences. While the current model works at the level of tracks and calorimeter clusters, our proposed approach scales naturally to cases where one considers the raw detector hits directly as an input, possibly allowing direct event reconstruction without having to tune clustering or tracking algorithms. It may be interesting to extend the hypergraph construction over dynamically binned events using the LSH approach. It may also be useful to construct features using semi-supervised or unsupervised learning from real data, to reduce the reliance on simulated datasets for supervised learning [66]. Furthermore, large-context models are continuously improving, and it may be interesting to apply the latest developments such as FlashAttention [67, 68] on the models in this paper. For improving throughput on e.g. CPUs, quantization has shown promise, and it may be interesting to investigate if this can be repeated for the type of models proposed for PF reconstruction. Finally, it is important to integrate the proposed machine learning (ML)-based reconstruction models into reconstruction frameworks such as CMSSW and Key4HEP.
Data and software availability
@@
The datasets are available in [49], the results including the trained model weight files in [69], and the software used for analysis in [70].

### 08 Mar 2024 v5

@@ 4 Conclusions
We have used a realistic simulation for the CLIC detector, conceptually similar to existing and future detector designs, to develop scalable machine learning models for full event particle flow reconstruction. We compare two scalable machine learning (ML) models: a graph neural network (GNN) model that uses locality sensitive hashing (LSH), and a kernel-based transformer using large-scale hyperparameter optimization (HPO). Both of these models avoid quadratic scaling in memory and computation, thus allowing the processing of events from highly granular detectors. We find that a GNN model significantly outperforms the kernel-based transformer alternative and the baseline Pandora-based particle-flow (PF) on the basis of individual particles as well as event-level quantities such as jets and total 3-momentum. Improved particle-level and event-level reconstruction can result in significant improvements for future flagship analyses such as those involving the Higgs boson decaying to bottom quarks. At the same time, the flexible and learnable LSH approach allows one to process complex events expected at future detectors with up to $10^{4}$ particles per event efficiently, while supporting various existing hardware accelerators such as Nvidia, AMD and Habana without additional porting effort. Our work contributes to the existing body of research by proposing a new, challenging open dataset for particle flow reconstruction studies, by defining relevant benchmarks and by identifying efficient models that can solve these benchmarks better than existing rule-based algorithms without additional tuning.
@@
Further research is possible in several directions. First, it would be useful to repeat this exercise using the simulation from detectors that are taking data in Run 3 of the Large Hadron Collider (LHC), e.g. in CMS, to study the performance of the reconstruction in more realistic conditions, and also study the reconstruction performance on real data. Second, we are currently using a simple particle-based loss, while the use of contrastive-adversarial learning methods may allow one to account for event-level discrepancies more effectively [65]. Third, the hypergraph model [21] shows promising physics performance but currently only supports small input sequences. It may be interesting to extend the hypergraph construction over dynamically binned events using the LSH approach. While the current model works at the level of tracks and calorimeter clusters, our proposed approach scales naturally to cases where one considers the raw detector hits directly as an input, possibly allowing direct event reconstruction without having to tune clustering or tracking algorithms. It may also be useful to construct features using semi-supervised or unsupervised learning from real data, to reduce the reliance on simulated datasets for supervised learning [66]. Furthermore, large-context models are continuously improving, and it may be interesting to apply the latest developments such as FlashAttention [67, 68] on the models in this paper. For improving throughput on e.g. CPUs, quantization has shown promise, and it may be interesting to investigate if this can be repeated for the type of models proposed for PF reconstruction. Finally, it is important to integrate the proposed ML-based reconstruction models into reconstruction frameworks such as CMSSW and Key4HEP.
Data availability
@@
Our datasets are published following the findable, accessible, interoperable, and reusable principles. The datasets are available in [49], the results including the trained model weight files in [69]
Code availability
Our code is published following the findable, accessible, interoperable, and reusable principles. The code used for analysis is available in [70].